Fermi National Laboratory
 

Accelerator Update for May 28, 2007 - May 30, 2007

 

Monday May 28

The day shift began with Operations monitoring the Recycler stash (206.1E10), with the Antiproton Source (Pbar) stacking (22.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the next store.

Operations put the Tevatron (TeV) into shot setup at 8:07 AM.

During the TeV tune up, Operations reported at 9:29 AM that the TeV was aborting. Operators investigated and discovered a Lambertson magnet power supply (T:ILAM) off.

Operations established store 5465 at 10:45 AM with an initial luminosity of 181.19E30.

Operations reported at 3:56 PM that the Recycler had stashed antiprotons twice during the day shift.

At 8:40 PM, the MI power supply interlock loop tripped off. Operators investigated and found that the VCB at MI-10L had opened.

At 9:04 PM, Operators reported that they had tried to close the MI-10L VCB, but it would not stay closed. Operators called in two power supply experts.

At 11:45 PM, the power supply experts reported that the MI-10L VCB was mechanically sound. The problem was due to a local interrupt on the interlock bypass loop. They continued to investigate.

Operations reported at 11:54 PM that the Recycler had stashed twice during the evening shift.

Tuesday May 29

The midnight shift began with Operations monitoring store 5465 (54.51E30) and the stash (141.7E10), with Pbar stacking (30.4E10), with NuMI waiting on MI beam, and with MiniBooNE taking beam.

Power supply experts working on the MI-10L VCB contacted another expert who suggested at 12:41 AM that they disable a certain relay and then try to close the VCB. Later, this expert came in to help diagnose the problem.

At 2:17 AM, experts reported that they could not find the problem with MI-10L and bypassed the system. They brought the MI-20L power supply on line.

Operations established beam through MI at 2:24 AM. Operations resumed sending beam to Pbar and NuMI.

Operations reported at 7:57 AM that the Recycler had stashed twice during the midnight shift.

At 7:58 AM, the TeV sector B3 MADC failed. An expert investigated and reported that it was beyond help. He will replace it when the current store ends.

Operations terminated store 5465 at 10 AM.

Controls experts replaced the TeV B3 MADC at 10:36 AM. The sector B3 readbacks resumed.

Operations made CDF ready at 10:13 AM for a 40-minute access.

Operations put the TeV through a dry squeeze at 11:06 AM.

Operations put the TeV into shot setup at 11:56 AM.

At 12:13 PM, Operations reported that they were having trouble getting beam through MI. Operators investigated and discovered that the MiniBooNE beam was again causing the problem. Operations lowered the MiniBooNE rep rate at 12:18 PM and this fixed the problem.

Operations established store 5467 at 3:28 PM with an initial luminosity of 196.64E30.

Operations reported at 3:35 PM that the MI-10 power supply VCB problems were resolved. They replaced a bad relay.

At 3:56 PM, the Pbar lithium lens power supply tripped off. After several tries, Operations reset the supply.

Operations reported at 4 PM that the Recycler had stashed once during the day shift.

Operations reported at 4:14 PM that Pbar experts were conducting Debuncher orbit measurements.

TeV experts began a sextupole study at 5:09 PM. CDF and D0 put their detectors into standby for this study, which took about 30-minutes.

Pbar experts ended their measurements and resumed stacking around 5:15 PM.

At 5:24 PM, the Pelletron tripped off. Experts investigated and discovered that its power supply current had drifted too high. Their investigations continued.

TeV experts completed their sextupole studies at 6:20 PM.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Wednesday May 30

The midnight shift began with Operations monitoring store 5467 (73.39E30) and the stash (119.5E10), with Pbar stacking (32.8E10), and with NuMI and MiniBooNE taking beam.

At 12:35 AM, the MI-8 transport line suffered a vacuum burst. Operators recovered the line.

At 2:52 AM, the NuMI beam permit tripped off due to a bad beam position. It reset.

At 2:53 AM, Operations reported that a poor Booster tune was causing MI problems. Operators turned off the Booster beam, set up plots to track what was happening, and then turned the beam back on. Unfortunately, the problem went away; the tune was fine and so were the beam positions.

Operations reported at 7:30 AM that Pbar was stacking better without using the stacking aggregate. Operators had been monitoring the stacktail cooling and adjusting it manually when it was needed.

Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow CDF an access when store 5467 is terminated later today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/30/2007    email Fermilab